# RAG-R1 : Incentivize the Search and Reasoning Capabilities of LLMs through Multi-query Parallelism

### v4

@@ 4 Multi-query Parallelism
By enabling the model to adaptively leveraging internal and external knowledge during reasoning, the capability boundaries of the model are significantly extended. However, existing methods generate only a single search query whenever external retrieval is required, which results in substantial retrieval iterations and limited knowledge acquired from retrieval. To enhance inference efficiency and the model’s capabilities, we further expand the generation and retrieval process within the framework from single-query mode to multi-query parallelism.
Specifically, we guide the model to generate multiple search queries whenever external retrieval is required. The external retrieval system then performs parallel searches with these queries and returns the results to the model in a JSON format, ensuring clear alignment between the search queries and the retrieved documents. Compared to single-query retrieval, multi-query parallel retrieval decreases the total number of retrievals needed, while maintaining comparable time for each retrieval. This approach is particularly beneficial for comparison-type questions, allowing for improved inference efficiency without compromising the model’s effectiveness. Moreover, by generating multiple search queries for parallel retrieval, the model can obtain more comprehensive and diverse information in a single interaction, which aids in subsequent decision-making. This enriched information acquisition further broadens the model’s capability boundaries, thereby augmenting its overall effectiveness.
Table 2 illustrates the system instructions for generating multiple search queries. We restrict the model to generate at most three parallel search queries simultaneously. The retrieval system returns the results in a JSON format which contains:
@@
documents: A list containing the retrieved documents corresponding to the search queries.
Utilizing the above format allows the model to recognize the alignment between search queries and retrieved documents. The remaining aspects of the training process are consistent with those outlined in Section 3. More case studies are available in Appendix B
@@ 5.1 Experimental Settings
Datasets and Evaluation Metrics We evaluate our models on seven benchmark datasets: (1) General Question Answering: NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and PopQA (Mallen et al., 2023). (2) Multi-Hop Question Answering: HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020), Musique (Trivedi et al., 2022), and Bamboogle (Press et al., 2023). HotpotQA is in-domain benchmark since a portion of its training sets is used for training while others serve as out-of-domain benchmarks to assess our model’s generalization capabilities. For evaluation metric, we utilize Exact Match (EM) following Yu et al. (2024).
Baselines We compare our method against the following baselines: (1) Naive Generation: Generating answers directly without retrieval. (2) Standard RAG: Traditional retrieval-augmented generation systems. (3)RAG-CoT Methods: Integration of retrieval-augmented generation with chain-of-thought prompts, such as IRCoT (Trivedi et al., 2023) and Search-o1 (Li et al., 2025). (4) RAG-RL Methods: Utilizing reinforcement learning to allow the model to autonomously perform retrieval during inference, such as Search-R1 (Jin et al., 2025) and R1-Searcher (Song et al., 2025).
Implementation Details We utilize Qwen-2.5-7B-Instruct as the backbone models for our training and all baselines. For retrieval, we use the English Wikipedia provided by KILT (Petroni et al., 2021) in 2019 as retrieval corpus, segmented into 100-word passages with appended titles, totaling 29 million passages. We employ BGE-large-en-v1.5 (Chen et al., 2023) as the text retriever and set the number of retrieved passages to 3 across all retrieval-based methods to ensure fair comparison. The maximum number of retrieval is set to 4. More details on the experimental settings are provided in Appendix A.
@@ 5.2 Main results
The main results comparing RAG-R1 with baseline methods across the seven datasets are presented in Table 3. RAG-R1-sq denotes our method operating in single-query mode while RAG-R1-mq represent the overall framework incorporating multi-query parallelism. From the results, we can obtain the following key observations:
Significant improvements across all datasets Our method achieves significant improvements compared to all baseline methods across both General QA and mutil-hop QA benchmarks, including both CoT-based methods and RL-based methods. Specifically, RAG-R1-mq outperforms the best RL-based method R1-Searcher by 13.2% across all datasets. These results demonstrate that our approach enables the model to effectively utilize both internal and external knowledge throughout the reasoning process.
Effectiveness of Multi-query Parallelism RAG-R1-mq outperforms all RL-based approaches that generate a single search query whenever external retrieval is required, including ours own RAG-R1-sq. This highlights that multi-query parallelism not only boosts inference efficiency but also effectively utilizes comprehensive and diverse information for reasoning, thereby improving the model’s performance. Further details are available in Section 6.2.
Excellent generalization Ability Despite utilizing only a subset of the HotpotQA training data, our models achieve significant improvements on in-domain datasets and exhibit excellent generalization capabilities across out-of-domain datasets, such as 2WikiMultiHopQA and Musique. This demonstrates that our models have effectively learned to reason and retrieve information for diverse questions, which proves the effectiveness of our approach across various scenarios requiring retrieval. Moreover, it can effectively extend to online search, as detailed in Section 6.3.
@@
| Methods | HotpotQA | 2Wiki | Avg. | |||
| Time | RC | Time | RC | Time | RC | |
| Search-R1 | 7.79 | 2.44 | 8.90 | 3.01 | 8.35 | 2.73 |
| R1-Searcher | 10.98 | 2.31 | 10.93 | 2.40 | 10.96 | 2.36 |
| RAG-R1-sq | 7.69 | 2.13 | 8.72 | 2.93 | 8.21 | 2.53 |
| RAG-R1-mq | 6.72 | 1.87 | 8.11 | 2.42 | 7.42 | 2.15 |
@@ 6.1 Ablation Study
SFT and RL As shown in Table 4, w/o SFT removes the initial format learning SFT while w/o RL removes the entire RL training stage. The results demonstrate the necessity and effectiveness of both SFT and RL in our training framework, which together improve the model’s performance. Specifically, w/o SFT struggles to leverage internal and external knowledge, resulting in decreased performance. In contrast, w/o RL restricts the model’s capability to correctly answer questions through reasoning. The considerable improvement achieved through RL highlight its ability to significantly enhance the model’s capability.
RL Data Selection We further validate the effectiveness of data selection during the RL process. We trained the models using all samples with incorrect answers and 25% of the samples with correct answers in Section 3.1.1, denoted as RAG-R1-sq w/o Filter and RAG-R1-mq w/o Filter respectively. The decline in performance indicates that careful data selection plays a crucial role in enhancing the effectiveness of RL training. Unanswerable samples do not facilitate the model’s improvement and may even detrimentally impact the training process.
@@ 6.2 Effects of Multi-query Parallelism
As shown in Table 5, we record the average inference time (in seconds) and average retrieval count for different methods on HotpotQA and 2WikiMultiHopQA using A100 GPUs without employing any inference acceleration technique. The results show that RAG-R1-mq achieves the lowest inference time and retrieval count, indicating that multi-query parallelism can significantly enhance the efficiency of the model. Furthermore, according to Table 3, our approach incorporating multi-query parallelism achieves the highest accuracy, demonstrating its effectiveness in enhancing the model’s capabilities. By comparing the responses of single-query mode and multi-query parallelism, we find that the improvement in effectiveness mainly comes from two aspects: (1) Multi-query parallelism reduces the number of retrieval while obtaining the correct answer, whereas single-query often fails to find the answer due to the limitation on the maximum number of retrieval. (2) Multi-query parallelism is capable of rewriting the query based on all existing information and conducting an additional retrieval when no effective information is found, whereas single-query mode lacks this ability. More case studies can be found in Appendix B. Consequently, multi-query parallelism not only boosts the model’s inference efficiency but also expands the boundaries of its capabilities.
@@ 6.3 Generalization to Online Search
Considering training efficiency and cost, we employ a local dense embedding retrieval system utilizing Wikipedia as a static external retrieval corpus throughout the training process. This contrasts with most real-world applications, which depend on online web search. To demonstrate the generalization capability of RAG-R1 in such scenarios, we assess our models’ performance on two benchmarks: Bamboogle and a randomly selected set of 500 samples from HotpotQA, utilizing online web search—a setting not encountered during training. Specifically, during inference, whenever retrieval is necessary, we leverage the Google API to perform real-time web searches and obtain relevant web pages. Subsequently, we utilize BeautifulSoup4 to scrape information from these pages and employ GPT-4o-mini to generate concise summaries, which are then incorporated into the reasoning process. As illustrated in Figure 3, RAG-R1-mq achieves the best EM score among all compared methods, highlighting its strong adaptability to online search scenarios. These findings suggest that our approach equips the model to dynamically retrieve information during the reasoning process, rather than merely memorizing response formats.
@@ 7 Conclusion
In this paper, we propose RAG-R1, a novel training framework that enables LLMs to adaptively leverage internal and external knowledge during the reasoning process and improves the reasoning ability to answer questions correctly. We further expand the generation and retrieval processes within the framework from single-query mode to multi-query parallelism to reduce the model’s inference time and enhance its capabilities. Extensive experiments on seven question answering benchmarks demonstrate the effectiveness of our method. In particular, our method utilizing multiple-query parallelism outperforms the strongest baseline by up to 13.2% and decreases inference time by 11.1%. Additionally, we conduct a comprehensive ablation analysis of its key design elements to assess the effectiveness of each component.
@@ Appendix A Implementation Details
We utilize Qwen-2.5-7B-Instruct as the backbone models for our training and all baseline models. For retrieval, we use the English Wikipedia provided by KILT (Petroni et al., 2021) in 2019 as retrieval corpus, segmented into 100-word passages with appended titles, totaling 29 million passages. We employ BGE-large-en-v1.5 (Chen et al., 2023) as the text retriever and set the number of retrieved passages to 3 across all retrieval-based methods to ensure fair comparison. The maximum number of retrieval iterations is set to 5.
@@
In Retrieval-Augmented RL, we collect 5022 samples for single-query training and 6015 samples for multi-query parallelism training. We split 95% of all the samples into a training set and used the remaining as a validation set. For the PPO variant, we set the learning rate of the policy LLM to 1e-6 and that of the value LLM to 1e-5. The training step is 500, with warm-up ratios of 0.285 and 0.015 for the policy and value models, respectively. We use Generalized Advantage Estimation (GAE) (Schulman et al., 2018) with parameters and . All the trainings are conducted on a single node with 8 A100 GPUs. We configure the batch settings as follows: a total batch size of 512, a mini-batch size of 128, and a micro-batch size of 32. The maximum sequence length is 8,192 tokens, with a maximum response length of 500 and retrieved content lengths set to a maximum of 1000 tokens for each search query.
For evaluation, we set the maximum number of retrieval iterations to 5, with a maximum response length of 1024 for each round and the temperature setting of 0.7.
Appendix B Case Study

### v5

@@ 4 Multi-query Parallelism
Enabling the model to adaptively leverage internal and external knowledge during reasoning significantly extends its capability boundaries. However, existing methods are constrained to a single-query approach, which presents two fundamental weaknesses: prohibitive latency from serial retrieval iterations, and inherent brittleness due to knowledge confinement. To overcome these limitations, we replace this fragile, sequential process with multi-query parallelism. This architectural shift is designed not merely to enhance capabilities, but to directly tackle the root causes of failure—it reduces latency by parallelizing retrieval and fortifies robustness by gathering diverse evidence simultaneously.
Specifically, we guide the model to generate multiple search queries whenever external retrieval is required. The external retrieval system then performs parallel searches with these queries and returns the results to the model in a JSON format, ensuring clear alignment between the search queries and the retrieved documents. The adoption of multi-query parallelism directly confronts the two fundamental weaknesses of the single-query approach: prohibitive latency and inherent brittleness. First, by executing retrievals in parallel, it dismantles the serial dependency that causes prohibitive latency, leading to a substantial reduction in overall inference time. Second, it shatters the knowledge confinement of a single query by gathering diverse evidence simultaneously. This enriched information base fortifies the model’s reasoning process, dramatically enhancing its robustness against suboptimal queries and unrecoverable paths. This dual improvement is particularly salient for complex tasks like comparison-type questions, where synthesizing multiple perspectives is crucial for accuracy.
Table 2 illustrates the system instructions for generating multiple search queries. We restrict the model to generate at most three parallel search queries simultaneously. The retrieval system reorganized the search results in a JSON format which contains:
@@
documents: A list containing the at most three retrieved documents corresponding to the search queries.
Utilizing the above format allows the model to recognize the alignment between difference search queries and corresponding retrieved documents. The remaining aspects of the training process are consistent with those outlined in Section 3. Detailed cases of multi-query methods are presented in the Appendix C, showcasing its ability and efficiency to handle complex and multi-hop queries.
@@ 5.1 Experimental Settings
Datasets and Evaluation Metrics. We evaluate our models on seven benchmark datasets: (1) General Question Answering: NQ (Natural_Questions), TriviaQA (JoshiTriviaQA2017), and PopQA (mallen2023trustlanguagemodelsinvestigating). (2) Multi-Hop Question Answering: HotpotQA (yang2018hotpotqa), 2WikiMultiHopQA (xanh2020_2wikimultihop), Musique (trivedi2021musique), and Bamboogle (press-etal-2023-measuring). HotpotQA is in-domain benchmark since a portion of its training sets is used for training while others serve as out-of-domain benchmarks to assess our model’s generalization capabilities. For evaluation metric, we utilize Exact Match (EM) score following rankrag and Retrieval Count (RC).
Baselines. We compare our method against the following baselines: (1) Direct Inference: Generate answers based on the model’s internal knowledge. (2) Standard RAG: Traditional retrieval-augmented generation systems. (3)RAG-CoT Methods: Integration of retrieval-augmented generation with chain-of-thought prompts, such as IRCoT (IRCoT) and Search-o1 (Search-o1). (4) RAG-RL Methods: Utilizing reinforcement learning to allow the model to autonomously perform retrieval during inference, such as Search-R1 (Search-R1) and R1-Searcher (R1-Searcher).
Implementation Details. We adopt Qwen-2.5-7B-Instruct as the base model for both our method and all baseline approaches, ensuring a fair comparison under the same architectural backbone. For retrieval, we use the English Wikipedia provided by KILT (KILT) as retrieval corpus, segmented into 100-word passages with appended titles, totaling 29 million passages. We employ BGE-large-en-v1.5 (bge_m3) as the text retriever and set the number of retrieved passages to 3 across all retrieval-based methods to ensure fair comparison. For evaluation, the number of retrieval iterations was left uncapped.
In Format Learning SFT, we collect 18994 samples for single-query training and 19303 samples for multi-query parallelism training. We perform full-parameter SFT for 5 epochs and chose the checkpoint with the lowest validation loss for evaluation and subsequent RL training.
In Retrieval-Augmented RL, we collect 5022 samples for single-query training and 6015 samples for multi-query parallelism training. We split 95% of all the samples into a training set and used the remaining as a validation set. For the PPO variant, we set the learning rate of the policy LLM to 1e-6 and that of the value LLM to 1e-5. The training step is 500, with warm-up ratios of 0.285 and 0.015 for the policy and value models, respectively. We use Generalized Advantage Estimation (GAE) (GAE) with parameters and . All the trainings are conducted on a single node with 8 A100 GPUs.
@@ 5.2 Main results
The main results comparing RAG-R1 with baseline methods across the seven datasets are presented in Table 3. RAG-R1-sq denotes our method operating in single-query mode while RAG-R1-mq represents the overall framework incorporating multi-query parallelism. From the results, we can obtain the following key observations:
Significant performance improvements across all datasets. Our method achieves significant improvements compared to all baseline methods across both General QA and mutil-hop QA benchmarks, including both CoT-based methods and RL-based methods. Specifically, RAG-R1-mq outperforms the best RL-based method, R1-Searcher, by 13.7% across all datasets. These results demonstrate that our approach enables the model to effectively utilize both internal and external knowledge throughout the reasoning process.
Effectiveness of Multi-query Parallelism. RAG-R1-mq consistently outperforms all single-query RL-based methods, notably achieving a 1.8% average gain over our own single-query counterpart, RAG-R1-sq. This result reveals a key insight: multi-query parallelism is not merely an efficiency optimization, but a direct remedy for the inherent brittleness of single-query methods. By shattering knowledge confinement, it empowers the model to synthesize diverse evidence, fundamentally bolstering its reasoning robustness against failures. Further details are available in Section 6.2.
Excellent generalization Ability. Despite utilizing only a subset of the HotpotQA training data, our models achieve significant improvements on in-domain datasets and exhibit excellent generalization capabilities across out-of-domain datasets, such as 2WikiMultiHopQA and Musique. This demonstrates that our models have effectively learned to reason and retrieve information for diverse questions, which proves the effectiveness of our approach across various scenarios requiring retrieval. Moreover, it can effectively extend to online search, as detailed in Section 6.3.
@@ 6.1 Ablation Study
SFT and RL As shown in Table 4, w/o SFT removes the initial format learning SFT while w/o RL removes the entire RL training stage. The results demonstrate the necessity and effectiveness of both SFT and RL in our training framework, which together improve the model’s performance. Specifically, w/o SFT struggles to leverage internal and external knowledge and respond in a think-then-search format, resulting in decreased performance. In contrast, w/o RL restricts the model’s capability to correctly answer questions through reasoning. The considerable improvement achieved through RL highlight its ability to significantly enhance the model’s capability.
RL Data Selection We further validate the effectiveness of data selection during the RL process. We trained the models using all samples with incorrect answers and 25% of the samples with correct answers in Section 3.1, denoted as RAG-R1-sq w/o Filter and RAG-R1-mq w/o Filter respectively. The decline in performance indicates that careful data selection plays a crucial role in enhancing the effectiveness of RL training. Unanswerable samples do not facilitate the model’s improvement and may even detrimentally impact the training process.
| Methods | HotpotQA | 2Wiki | Avg. | |||
| Time | RI | Time | RI | Time | RI | |
| Search-R1 | 7.79 | 2.44 | 8.90 | 3.01 | 8.35 | 2.73 |
| R1-Searcher | 10.98 | 2.31 | 10.93 | 2.40 | 10.96 | 2.36 |
| RAG-R1-sq | 7.69 | 2.14 | 8.72 | 2.72 | 8.21 | 2.43 |
| RAG-R1-mq | 6.72 | 1.89 | 8.11 | 2.43 | 7.42 | 2.16 |
@@ 6.2 Effects of Multi-query Parallelism
As shown in Table 5, we record the average inference time (in seconds) and average retrieval iterations for different methods on HotpotQA and 2WikiMultiHopQA using A100 GPUs without employing any inference acceleration technique. The results show that RAG-R1-mq achieves the lowest inference time and requires the fewest retrieval iterations, indicating that multi-query parallelism can significantly enhance the efficiency of the model. As shown in Table 3, our multi-query approach (RAG-R1-mq) confirms its superior effectiveness. A detailed analysis against our single-query baseline (RAG-R1-sq) reveals this performance gain stems from two fundamental advantages: (1) More Efficient and Decisive Exploration: The MQ model’s parallel retrieval acts as a wide-ranging initial search, quickly identifying promising information pathways. This prevents the fruitless retrieval loops or premature terminations that plague the SQ model, which often lead to unrecoverable errors. (2) Adaptive Reasoning and Self-Correction: When initial retrievals fail, the MQ model demonstrates a crucial capability for self-correction. It synthesizes the complete context from all parallel queries to intelligently formulate a more informed new query. In contrast, the SQ model, locked into its linear path, lacks this adaptive mechanism, rendering it brittle to an initially suboptimal query. More case studies can be found in Appendix C. In essence, multi-query parallelism fosters a more robust and adaptive reasoning process, expanding the model’s problem-solving capabilities beyond mere efficiency gains.
@@ 6.3 Generalization to Online Search
Considering training efficiency and cost, we employ a local dense embedding retrieval system utilizing Wikipedia as a static external retrieval corpus throughout the training process. This contrasts with most real-world applications, which depend on online web search. To demonstrate the generalization capability of RAG-R1 within this scenario, we assess our models’ performance on two benchmarks: Bamboogle and a randomly selected set of 500 samples from HotpotQA, utilizing online web search—a setting not encountered during training.
Specifically, during inference, whenever retrieval is necessary, we leverage the Google API to perform real-time web searches and obtain relevant web pages. Subsequently, we utilize BeautifulSoup4 to scrape information from these pages and employ GPT-4o-mini to generate concise summaries, which are then incorporated into the reasoning process. As illustrated in Figure 3, RAG-R1-mq achieves the best EM score among all compared methods, highlighting its strong adaptability to online search scenario. These findings suggest that our approach equips the model to dynamically retrieve information during the reasoning process, rather than merely memorizing response formats.
@@ 7 Conclusion
In this paper, we introduced RAG-R1, a novel training framework that enables LLMs to adaptively leverage internal and external knowledge, significantly enhancing their reasoning capabilities. The cornerstone of our work is the integration of multi-query parallelism, an architectural innovation designed to directly address the prohibitive latency and inherent brittleness of conventional single-query methods, thereby bolstering reasoning robustness and reducing inference time. Extensive experiments on seven QA benchmarks demonstrate the effectiveness of our method, which outperforms the strongest baseline by up to 13.7% and decreases inference time by 11.1%. This dual advancement confirms that our method achieves a superior trade-off, simultaneously boosting the model’s reasoning robustness and inference efficiency.
@@ Appendix A Implementation Details
We adopt Qwen-2.5-7B-Instruct as the base model for both our method and all baseline approaches, ensuring a fair comparison under the same architectural backbone. For retrieval, we use the English Wikipedia provided by KILT (KILT) as retrieval corpus, segmented into 100-word passages with appended titles, totaling 29 million passages. We employ BGE-large-en-v1.5 (bge_m3) as the text retriever and set the number of retrieved passages to 3 across all retrieval-based methods to ensure fair comparison.
@@
In Retrieval-Augmented RL, we collect 5022 samples for single-query training and 6015 samples for multi-query parallelism training. We split 95% of all the samples into a training set and used the remaining as a validation set. For the PPO variant, we set the learning rate of the policy LLM to 1e-6 and that of the value LLM to 1e-5. The training step is 500, with warm-up ratios of 0.285 and 0.015 for the policy and value models, respectively. We use Generalized Advantage Estimation (GAE) (GAE) with parameters and . All the trainings are conducted on a single node with 8 A100 GPUs. We configure the batch settings as follows: a total batch size of 512, a mini-batch size of 128, and a micro-batch size of 32. The maximum sequence length is 8,192 tokens, with a maximum response length of 500 and retrieved content lengths set to a maximum of 1000 tokens for each search query.
For evaluation, the number of retrieval iterations was left uncapped, with a maximum response length of 1024 for each round and the temperature setting of 0.7.
Appendix B SFT Data Segmentation
| Raw Data: |
| Question: Are the magazines, Nat Mags and Plaza Kvinna published in the same country? |
| thinkTo determine if the magazines Nat Mags and Plaza Kvinna are published in the same country, I need to find out the countries of publication for both magazines./think |
| search Nat Mags country of publication, Plaza Kvinna country of publication /search |
| information |
| { ”query”: |
| ”Nat Mags country of publication”, |
| ”Plaza Kvinna country of publication” , |
| ”documents”: |
| ”Doc 1: Nat Mags Nat Mags Nat Mags (short for National Magazine Company) was a British magazine publisher based in London…”, |
| ”Doc 1: Plaza Kvinna Plaza Kvinna Plaza Kvinna (meaning P̈laza Womanïn English) is a Swedish language monthly women’s magazine published in Stockholm, Sweden…” |
| /information |
| thinkFrom the search results, Nat Mags is a British magazine publisher based in London, indicating that the magazines under Nat Mags are published in the United Kingdom. Plaza Kvinna, on the other hand, is a Swedish language monthly women’s magazine published in Stockholm, Sweden. Therefore, these two magazines are not published in the same country./think |
| answer No /answer |
| SFT Data after Segmentation: |
| Data 1: |
| input: Question: Are the magazines, Nat Mags and Plaza Kvinna published in the same country? |
| output: thinkTo determine if the magazines Nat Mags and Plaza Kvinna are published in the same country, I need to find out the countries of publication for both magazines./think |
| search Nat Mags country of publication, Plaza Kvinna country of publication /search |
| Data 2: |
| input: Question: Are the magazines, Nat Mags and Plaza Kvinna published in the same country? |
| thinkTo determine if the magazines Nat Mags and Plaza Kvinna are published in the same country, I need to find out the countries of publication for both magazines./think |
| search Nat Mags country of publication, Plaza Kvinna country of publication /search |
| information |
| { ”query”: |
| ”Nat Mags country of publication”, |
| ”Plaza Kvinna country of publication” , |
| ”documents”: |
| ”Doc 1: Nat Mags Nat Mags Nat Mags (short for National Magazine Company) was a British magazine publisher based in London…”, |
| ”Doc 1: Plaza Kvinna Plaza Kvinna Plaza Kvinna (meaning P̈laza Womanïn English) is a Swedish language monthly women’s magazine published in Stockholm, Sweden…” |
| /information |
| output: thinkFrom the search results, Nat Mags is a British magazine publisher based in London, indicating that the magazines under Nat Mags are published in the United Kingdom. Plaza Kvinna, on the other hand, is a Swedish language monthly women’s magazine published in Stockholm, Sweden. Therefore, these two magazines are not published in the same country./think |
| answer No /answer |
Appendix C Case Study